Circular No. 8829

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

GW LIBRAE E. O. Waagen, AAVSO, reports that this apparent WZ Sge-type cata-

clysmic variable is in outburst for the first time since its discovery in 1983 (cf. IAUC 3854). Selected visual magnitude estimates (see also CBET 922): Apr. 12.126 UT, [12.0 (P. Schmeer, Bischmisheim, Germany); 12.494, 13.8 (R. Stubbings, Drouin, Vic., Australia); 12.569, 13.4 (A. Pearce, Nedlands, W. Australia; revision to CBET 922); 12.617, 12.5 (Pearce; revision to CBET 922); 12.659, 11.8 (Stubbings); 12.853, 9.9 (Pearce).

SUPERNOVAE 2007av–2007bd Several supernovae have been discovered on unfiltered CCD frames:

2007av and 2007ax by R. Arbour (cf. IAUC 8496); and the rest by W. Li (LOSS/KAIT; cf. IAUC 8822; with R. Mostardi for 2007aw and 2007ay, with D. Winslow for 2007az and 2007ba, with N. Joubert for 2007bb, and with R. Prasad for 2007bc and 2007bd). Discovery observations:

SN  2007 UT  α_{2000}  δ_{2000}  Mag.  Offset
2007av Mar. 20.84 10^{h}34^{m}43.^{s}17 +11^{o}11′38′′.3 15.5 5′′.9 E, 14′′.9 S
2007aw Mar. 22.26 9 57 24.16 −19 21 23.1 18.1 3′′.3 E, 5′′.2 S
2007ax Mar. 21.98 8 22 43.26 +22 33 16.9 17.2 2′′.6 W, 5′′.5 N
2007ay Mar. 22.21 8 17 14.85 + 1 12 06.9 17.1 12′′.7 W, 16′′.4 S
2007az Mar. 6.27 8 25 23.80 +69 54 29.4 18.9  —
2007ba Mar. 29.49 15 16 42.63 + 7 23 47.8 18.2 11′′.5 E, 0′′.9 S
2007bb Apr. 2.15 7 01 07.46 +51 15 57.3 17.2 22′′.3 E, 14′′.0 S
2007bc Apr. 4.32 11 19 14.57 +20 48 32.5 16.1 29′′.2 W, 16′′.5 S
2007bd Apr. 4.18 8 31 33.28 − 1 11 58.0 18.4 7′′.6 E, 5′′.1 N

Additional information is available on Electronic Telegrams: 2007av in NGC 3279, CBETs 901 and 903 (type-II, discovered close to explosion time); 2007aw in NGC 3072, CBETs 902 and 908 (type-Ic, discovered a few days after maximum); 2007ax in NGC 2577, CBETs 904, 907, and 908 (subluminous type-Ia, discovered several days before maximum); 2007ay in UGC 4310, CBETs 905 and 908 (type-II, discovered shortly after explosion time); 2007az, CBETs 872 and 909 (type-Ib, discovered shortly before maximum); 2007ba in UGC 9798, CBETs 911 and 918 (type-Ia, discovered well before maximum); 2007bb in UGC 3627, CBETs 912 and 916 (type-IIn, discovered just after explosion); 2007bc in UGC 6332, CBETs 913 and 915 (type-Ia, discovered several days before maximum); 2007bd in UGC 4455, CBETs 914 and 915 (type-Ia, discovered > 1 week before maximum).

2007 April 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green